OpenStudy (anonymous):

The perimeter of a rectangular lot is 56 meters. The length exceeds the width by 12 meters. Find the length of the lot. 14 m 28 m 15 m 20 m

OpenStudy (anonymous):

l=12+w perimeter=2(l+w)=56 2(l+l-12)=56 2l=40 l=20

OpenStudy (anonymous):

l=20m

OpenStudy (anonymous):

see here it is given that length = 12 + width that is l =12 + w now see here as we know that the perimeter of a rectangle = 2(l+w) we have 56 m = 2(12+w+w) 56 m = 2(12+2w) 28m = 12 + 2w 28 - 12m = 2w 16m = 2w 16m/2 =w 8m = width and l = 12 + w that is l = 12 + 8 = 20m hence l = 20m and b = 8 m